OpenStudy (he66666):

How do you solve the derivative of 1000(r)^(-1) or 1000(r^(-1))? I know how to solve simple ones but this one confuses me because there's a constant in the front

OpenStudy (anonymous):

treat the constant as that... a constant if you had f(x)=Ax^2 and A was a constant the derivative would be f'(x)=A2x..... because all a constant does is multiply the function my some factor A

OpenStudy (he66666):

oh I see. thanks nadeem!

OpenStudy (anonymous):

I see how this can be confusing because \[d/dx(c)=0, \] Where c is a constant but \[d/dx=(c*f(x))'=c*f'(x)\]

OpenStudy (anonymous):

I'm sorry made a typo this is what i meant \[d/dx(c)=0,\] and c is a constant \[d/dx(cf(x))=(c∗f(x))'=c*f'(x)\]

OpenStudy (anonymous):

I've been up all night study for an exam...lol

OpenStudy (he66666):

Thanks a bunch. Hope you do well on your exam!